OpenStudy (anonymous):

An equation is shown below: 4(x - 3) - 5(x + 1) = 3 Which statement shows a correct next step in solving the equation? (1 point) Question 18 options: 1) The equation can become 4x - 3 - 5x + 1 = 3 by applying the associative property of multiplication. 2) The equation can become 4x - 3 - 5x + 1 = 3 by applying the distributive property. 3) The equation can become 4x - 12 - 5x - 5 = 3 by applying the distributive property. 4) The equation can become 4x - 12 - 5x - 5 = 3 by applying the associative property of multiplication.

OpenStudy (anonymous):

4x -12 - 5x - 5= 3 -x -17 = 3 -x = 3+ 17 x = -20